Xachs Inc. is a small biotech company with 100 million shares trading at $10/share and debt outstanding of $ 250 million. The firm has a levered beta of 1.00 and a pre-tax cost of debt of 4.5%. The appropriate risk-free rate is 3.5%, the marginal tax rate is 40% and the equity risk premium is 5%.
Estimate the current cost of capital for the firm. Now assume that the firm plans to borrow $ 500 million and buy back stock. This will triple the default spread on the debt (both new and existing), estimate the new cost of capital for the firm after the recapitalization.
Now assume that the firm does buy back stock with the $ 500 million and pays $11/share. Estimate the value per share for the remaining shareholders in the company. (You can assume that it is a zero growth stock in perpetuity)
How would your answer to c have changed, if the firm had been able to keep all of its existing debt at the existing interest rate of 4.5% even after the recapitalization?
1
Expert's answer
2012-06-22T10:06:41-0400
Cost of capital = Risk free rate of return + Beta x (market rate of return- risk free rate of return)
=
= 3.5 + 1*5 = 8.5%
Now assume that the firm plans to borrow $ 500 million and buy back stock. This will triple the default
spread on the debt (both new and existing), estimate the new cost of capital for
the firm after the recapitalization.
The cost of capital will decrease.
The value per share for the remaining shareholders in the company will be
$11/(5% - 3.5%) = $733.33